could you show me how to set up these equations?? Thanks word problem. the sum of two numbers is 30. five times the smaller exceeds twice the larger by 17. find the numbers.
and A collection of nickels and quarters totals 3.40. the totale number of coins is 24. how many of each are there?

Answer by mananth(16946)

the sum of two numbers is 30. five times the smaller exceeds twice the larger by 17. find the numbers.
...
let one number be x
other will be 30-x
..
5x=2(30-x)+17
5x=60-2x+17
5x+2x=60+17
7x=77
/7
x= 11
other number = 19
and A collection of nickels and quarters totals 3.40. the totale number of coins is 24. how many of each are there?
total coins = 24
let x be nickels
24-x will be quarters
..
value 340 cents
..
5x+25(24-x)=340
5x+600-25x=340
-20x=340-600
-20x=-260
x=13 nickels
quarters = 11
...
11*25+5*13=340
...
m.ananth@hotmail.ca